# Proximal Policy Distillation

### v1

@@ Abstract
We introduce Proximal Policy Distillation (PPD), a novel policy distillation method that integrates student-driven distillation and Proximal Policy Optimization (PPO) to increase sample efficiency and to leverage the additional rewards that the student policy collects during distillation. To assess the efficacy of our method, we compare PPD with two common alternatives, student-distill and teacher-distill, over a wide range of reinforcement learning environments that include discrete actions and continuous control (ATARI, Mujoco, and Procgen). For each environment and method, we perform distillation to a set of target student neural networks that are smaller, identical (self-distillation), or larger than the teacher network. Our findings indicate that PPD improves sample efficiency and produces better student policies compared to typical policy distillation approaches. Moreover, PPD demonstrates greater robustness than alternative methods when distilling policies from imperfect demonstrations. The code for the paper is released as part of a new Python library built on top of stable-baselines3 to facilitate policy distillation: ‘sb3-distill’.
@@ 1 Introduction
Regarding implementation, KD typically involves adding a distillation loss to the problem-specific (dataset) loss function, to train a student network to reproduce the same output values as the teacher model. On the other hand, most policy distillation methods only use a distillation loss, training the student through supervised learning rather than in a reinforcement learning setting. However, this overlooks the valuable information that can be gained from the rewards that the student collects during distillation. Exploiting these rewards could accelerate the distillation process, potentially enable the student to outperform the teacher, and reduce the risk of overfitting to imperfect teachers.
We introduce Proximal Policy Distillation (PPD), a novel policy distillation method that combines student-driven distillation and Proximal Policy Optimization (PPO) (Schulman et al., 2017). PPD enhances PPO by incorporating a distillation loss to either perform traditional distillation, or to act as skills prior. Specifically, PPD enables the student to accelerate learning through distillation from a teacher, while potentially surpassing the teacher’s performance.
@@ 2 Proximal Policy Distillation
We build on the framework developed by Czarnecki et al. (2019) since in its general form it already allows for the inclusion of environment rewards during distillation. The general form of the distillation gradient in this framework is given by
@@
$\nabla_{\theta}L(\theta)=\mathbb{E}_{q_{\theta}}\left[\sum_{t=1}^{|\tau|}-\nabla_{\theta}\log\pi_{\theta}(a_{t}|\tau_{t})\widehat{R_{t}}+\nabla_{\theta}\ell(\pi_{\theta},V_{\pi_{\theta}},\tau_{t})\right]$ (1)
We focus in particular on the case where the student is used as the exploration policy $q_{\theta}$ to collect trajectories $\tau=\{s_{t},a_{t},r_{t},s_{t+1},a_{t+1},\dots,r_{t+N}\}$ by interacting with the environment. The use of student policies to collect trajectories during distillation generally results in superior policies compared to using the teacher. This is because using the teacher policy tends to overfit to the states it visits, neglecting a significant portion of the state space that the well-performing teacher overlooks (Czarnecki et al., 2019).
We then extend the formulation by incorporating elements from Proximal Policy Distillation (PPO) (Schulman et al., 2017). Specifically, we use the PPO loss to integrate environment rewards into the objective, we enable the reuse of samples from a rollout buffer to increase sample efficiency, via importance sampling, and we introduce proximality constraints to improve stability.
Proximal Policy Distillation works as follows: a student agent is trained similarly to standard PPO, by alternating between filling up a rollout buffer of trajectories through interaction with the task environment, using the current policy $\pi_{\theta_{k}}$, and performing policy updates. Policy updates are calculated by mini-batch gradient descent using samples from the rollout buffer for a fixed number of epochs, to solve the following (implicitly) constrained optimization problem at each iteration:
@@
$\theta_{k+1}=\arg\max_{\theta}\mathbb{E}_{s,a\sim\pi_{\theta_{k}}}\left[L_{\text{PPO}}(s,a,\theta)-{\color[rgb]{0,0,1}\lambda\,\text{KL}(\pi_{\text{teacher}}(s)\|\pi_{\theta}(s))\max\left(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{k}}(a|s)},1-\epsilon\right)}\right]$
[…]
@@
where the constraint is enforced by the hyperparameter $\epsilon$. $L_{\text{PPO}}$ is the PPO-clip loss, $\lambda$ is a hyperparameter that balances the relative strength of the PPO and distillation losses, $\pi_{\theta_{k}}$ is the policy from the previous PPO iteration that was used to collect the rollout buffer (hence, $\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{k}}(a|s)}$ is the importance sampling ratio), $\pi_{\theta_{k+1}}$ is the resulting policy at the end of the iteration, which will be used to collect the next rollout, and $\pi_{\text{teacher}}$ is the teacher policy that we wish to use to guide learning (e.g., for distillation or as skills prior). Note that we clip the distillation objective in the same way as the PPO loss; however, since KL-divergence is always non-negative, we only need to clip the positive branch of the loss. The critic is trained from scratch using a regression loss, as in PPO. The full algorithm is reported in Appendix A (algorithm 1).
@@ 3 Empirical Evaluation
Evaluation was performed over a wide range of challenging reinforcement learning environments involving discrete and continuous action spaces, continuous control, and out-of-distribution generalization. Namely, we tested all distillation methods on three suites of environments:
@@
Figure 1: Training curves of PPD and student-distill in the setting of distillation onto a larger student network, showing the average episodic returns collected during distillation, averaged over five random seeds and all environments in each of three suites ‘atari’, ‘mujoco’, and ‘procgen’. The final mean training performance of the teacher models is shown as a dashed line. Results for procgen are calculated over training levels. We observe that PPD students achieve higher rewards than student-distill ones, and are generally more sample efficient.
@@
Table 1: Performance of student models of three sizes (smaller, same, larger), trained using the three distillation methods (PPD, student-distill, and teacher-distill). Evaluation is performed in a test setting (which in the case of ‘procgen’ corresponds to using a different set of levels). Results are calculated as fraction of teacher score for each environment and random seed, and then averaged by geometric mean. We omit the ’x’ symbols in all but the last row to reduce clutter.
env teacher smaller same-size larger
score td sd ppd td sd ppd td sd ppd
atari 2.19 0.89 1.05 1.06 1.11 1.16 1.19 1.15 1.14 1.25
mujoco 4322.42 0.94 0.97 0.97 0.96 1.0 1.01 0.94 1.0 1.0
procgen 5.16 0.59 0.66 0.75 0.75 0.82 0.98 0.83 0.93 1.02
all 0.83x 0.94x 0.97x 0.99x 1.05x 1.11x 1.03x 1.07x 1.14x
@@
Table 2: Performance of student models, trained using the three distillation methods (PPD, student-distill, and teacher-distill) using ‘imperfect teachers’ that are artificially corrupted to decrease in performance. Results are calculated as a fraction of the original teacher score for each environment and random seed, and then averaged by geometric mean. Distillation is performed on four Atari and four Procgen environments, and onto larger student networks.
@@
Figure 2: Performance of student models trained using PPD with different values of $\lambda\in\left\{0.5,1,2,5\right\}$. Results are averaged over 5 random seeds. Distillation is performed on four Atari environments and onto larger student networks. Dashed lines indicate $100\%$ (black) and $90\%$ (gray) of the average performance of the teacher models. A circle on each curve marks the last training step where the student’s performance is below 90% of the teacher’s. All curves were collected during training, as in Figure 1. Values in parentheses indicate the relative final test performance of the students with respect to the teacher.
@@ 4 Related Work
Most current policy distillation methods overlook the rewards obtained by the student during the distillation process. Instead, they frame policy distillation solely as a supervised learning task. In this approach, a dataset of state observations and teacher outputs is collected using either the student or teacher policies. The data is then stored in a distillation dataset, from which training mini-batches are sampled.
While previous work (Czarnecki et al., 2019) has presented a general framework for policy distillation that allows the inclusion of environment rewards, the authors show that many types of current PD methods may have difficulty converging in that case.
@@
PD as a form of supervised learning can be easily adapted to transfer functions between different types of models, such as value-based teachers (e.g., trained with DQN) and policy-based students. This proves beneficial in scenarios like reincarnating reinforcement learning (Agarwal et al., 2022), allowing continued training with alternate RL methods after distillation. For instance, Green et al. (2019) showed that teachers trained using DQN can be distilled into “PPO students,” allowing for subsequent fine-tuning with Proximal Policy Optimization (PPO) (Green et al., 2019). However, unlike in our work, PPO was not used during distillation, and the trajectories for the distillation loss were collected in a teacher-driven way, with the teacher interacting with the environment instead of the student. This method closely resembles the ‘teacher_distill’ strategy in our baseline, which we demonstrated to be less effective than PPD.
@@
An alternative way to perform policy distillation is through imitation learning (IL), where demonstrations from teacher models provide specific actions as hard targets, rather than using the typical soft targets (e.g., output probabilities). For example, Zhuang et al. (2023) use DAgger (Dataset Aggregation) (Ross et al., 2011) to distill five behaviors for a robot quadruped (climb, leap, crawl, tilt, run) into a single control policy, by selecting the appropriate teacher for each part of a training obstacle course. The advantage of DAgger over naive Behavior Cloning stems from its ability to query teacher actions for new states encountered by the student, which is always possible in policy distillation, rather than relying solely on teacher-collected demonstrations.
Finally, closest to our work is the approach by Schmitt et al. (2018), which combines a policy gradient loss with a distillation loss and student-driven distillation. The authors also propose a method to dynamically update the relative weighting $\lambda$ of the policy gradient loss and distillation losses for multiple teachers, using Population Based Training (PBT) (Jaderberg et al., 2017). The work then applies distillation in a manner similar to reincarnating RL (Agarwal et al., 2022), where a smaller teacher is distilled into a larger student. However, distillation and learning are combined, instead of being performed sequentially. The main difference with Proximal Policy Distillation is that the PPD extends PPO, while the work by Schmitt et al. (2018) is based on IMPALA (Espeholt et al., 2018).
@@ 5 Discussion and Conclusions
Future work should focus on extending PPD to use teachers as skills priors, especially in the case of multitask policy distillation. To make that practical, and to further help PPD students to achieve higher performance than their teacher, it will be beneficial to implement a dynamic balancing in the trade-off between the PPO and individual teacher distillation losses, as for example done by Schmitt et al. (2018), who used Population Based Training to update the $\lambda_{k}$ parameters during training.

### v2

@@ Abstract
We introduce Proximal Policy Distillation (PPD), a novel policy distillation method that integrates student-driven distillation and Proximal Policy Optimization (PPO) to increase sample efficiency and to leverage the additional rewards that the student policy collects during distillation. To assess the efficacy of our method, we compare PPD with two common alternatives, student-distill and teacher-distill, over a wide range of reinforcement learning environments that include discrete actions and continuous control (ATARI, Mujoco, and Procgen). For each environment and method, we perform distillation to a set of target student neural networks that are smaller, identical (self-distillation), or larger than the teacher network. Our findings indicate that PPD improves sample efficiency and produces better student policies compared to typical policy distillation approaches. Moreover, PPD demonstrates greater robustness than alternative methods when distilling policies from imperfect demonstrations. The code for the paper is released as part of a new Python library built on top of stable-baselines3 to facilitate policy distillation: https://github.com/spiglerg/sb3_distill.
@@ 1 Introduction
Regarding implementation, KD typically involves adding a distillation loss to the problem-specific (dataset) loss function, to train a student network to reproduce the same output values as the teacher model. While several recent works have explored combining policy distillation with reinforcement learning objectives (Rajeswaran et al., 2017; Weihs et al., 2021; Nguyen et al., 2022; Shenfeld et al., 2023), many policy distillation methods still use only a distillation loss, training the student through supervised learning rather than in a reinforcement learning setting. However, this overlooks the valuable information that can be gained from the rewards that the student collects during distillation. Exploiting these rewards could accelerate the distillation process, potentially enable the student to outperform the teacher, and reduce the risk of overfitting to imperfect teachers.
We introduce Proximal Policy Distillation (PPD), a novel policy distillation method that combines student-driven distillation and Proximal Policy Optimization (PPO) (Schulman et al., 2017). PPD enhances PPO by incorporating a distillation loss to either perform traditional distillation, or to act as skills prior. Specifically, PPD enables the student to accelerate learning through distillation from a teacher, while potentially surpassing the teacher’s performance. While PPD is designed and evaluated primarily with fully-observable students, it can also be applied in partially-observable tasks, for instance by using a student policy with memory (e.g., LSTM-based), provided that the teacher policy has access to the full environment state.
@@ 2 Proximal Policy Distillation
We build on the framework developed by Czarnecki et al. (2019) since in its general form it already allows for the inclusion of environment rewards during distillation. The general form of the distillation gradient in this framework is
@@
$\nabla_{\theta}L(\theta)=\mathbb{E}_{q}\left[\sum_{t=1}^{|\tau|}\underbrace{-\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid\tau_{t})\widehat{R_{t}}}_{\text{\small policy-gradient term }}+\underbrace{\nabla_{\theta}\ell(\pi_{\theta},V_{\pi_{\theta}},\tau_{t})}_{\text{\small distillation term }}.\right]$ (1)
Where $\theta$ are the parameters of the student policy $\pi_{\theta}$ that we wish to train (possibly in conjunction with its value function $V_{\pi_{\theta}}$), and $\tau=\{s_{1},a_{1},r_{1},s_{2},a_{2},\dots,r_{|\tau|}\}$ are trajectories sampled by interacting with the target environment using a control policy $q$. $\widehat{R_{t}}$ is a reward signal for timestep $t$, which can take many forms as in vanilla Policy Gradient (e.g., a sum of episode returns $\sum_{t^{\prime}=1}^{|\tau|}r_{t}$, (discounted) sum of returns-to-go $\sum_{t^{\prime}=t}^{|\tau|}r_{t^{\prime}}$, the advantage function $A^{\pi}(s_{t},a_{t})$, etc). Finally, $\ell(\pi_{\theta},V_{\pi_{\theta}},\tau_{t})$ is an auxiliary term for the chosen distillation method (for example, KL-divergence between teacher and student policies).
We focus on the setting where the student itself acts as the control policy, i.e., $q=\pi_{\theta}$, since it typically yields better performance after distillation than using the teacher. In fact, relying on the teacher for exploration biases the data toward its limited state-visitation distribution, leaving many states unexplored (Czarnecki et al., 2019). In this paper, we refer to distillation methods that use the student as the control policy as student-driven, and those that use the teacher as teacher-driven.
To obtain Proximal Policy Distillation (PPD), we extend the formulation from Eq. 1 by replacing the vanilla policy-gradient term with the Proximal Policy Optimization (PPO) surrogate (Schulman et al., 2017). This requires three modifications. (i) We improve sample efficiency by extending the original formulation to re-use data from a rollout buffer for several epochs, instead of using the sampled trajectories only once per update. As such, the control policy used to collect data for the $k+1$ PPD update is the student policy from the previous step $q=\pi_{\theta_{k}}$, and we need to introduce importance-sampling weights $\rho_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{k}}(a_{t}\mid s_{t})}$. (ii) We introduce clipping to implement a proximality constraint to improve training stability, as in PPO-clip. (iii) We choose KL-divergence as distillation loss, which we also clip
@@
$\ell(\pi_{\theta},V_{\pi_{\theta}},\tau_{t})=\lambda\,\mathrm{KL}\!\left(\pi_{\text{teacher}}(\cdot\mid s_{t})\;\|\;\pi_{\theta}(\cdot\mid s_{t})\right)\max\!\bigl{(}\rho_{t}(\theta),1-\epsilon\bigr{)}$
Note that clipping the KL term is not required when using forward KL-divergence. However, simultaneous distillation from multiple teachers may benefit from using reverse KL-divergence, for which clipping is necessary. We thus keep the clipping term to retain a more general formulation.
Rather than write another gradient, we package these ingredients into an implicit optimization problem which is iteratively updated by mini-batch SGD:
$\theta_{k+1}=\arg\max_{\theta}\mathbb{E}_{s,a\sim\pi_{\theta_{k}}}\left[\;L_{\text{PPO}}(s,a,\theta)-{\color[rgb]{0,0,1}\lambda\,\text{KL}(\pi_{\text{teacher}}(s)\|\pi_{\theta}(s))\max\left(\frac{\pi_{\theta}(a\mid s)}{\pi_{\theta_{k}}(a\mid s)},1-\epsilon\right)}\;\right]$
$L_{\text{PPO}}(s,a,\theta)=\min\left(\frac{\pi_{\theta}(a\mid s)}{\pi_{\theta_{k}}(a\mid s)}\hat{A}(s,a),g(\epsilon,\hat{A}(s,a))\right),\hskip 21.68121ptg(\epsilon,A)=\begin{cases}(1+\epsilon)A,&A\geq 0\\ (1-\epsilon)A,&A<0\end{cases}$ (2)
@@
where the proximality constraint is enforced by the hyperparameter $\epsilon$. $L_{\text{PPO}}$ is the PPO-clip loss, $\lambda$ is a hyperparameter that balances the relative strength of the PPO and distillation losses. $\pi_{\theta_{k}}$ is the policy from the previous PPO iteration that was used to collect the rollout buffer, $\pi_{\theta_{k+1}}$ is the resulting policy at the end of the iteration, which will be used to collect the next rollout, and $\pi_{\text{teacher}}$ is the teacher policy that we wish to use to guide learning. The critic is trained from scratch using a regression loss, as in PPO.
The full algorithm is reported in Appendix A (algorithm 1).
@@ 3 Empirical Evaluation
Evaluation was performed on environments from Atari, Mujoco and Procgen because they span three axes that most affect performance in reinforcement learning and policy distillation: (i) state-space complexity – low-dimensional states (Mujoco) vs. high-dimensional pixel observations (Atari & Procgen); (ii) action spaces – discrete (Atari & Procgen) vs. continuous (Mujoco); (iii) out-of-distribution generalization – identical train/test environment (Atari & Mujoco) vs. procedurally generated train/test splits (Procgen). All tasks were episodic but discounted, thus differring from the strictly undiscounted formulation in (Czarnecki et al., 2019).
Namely, we used the following environments:
@@
Figure 1: Training curves of PPD and student-distill in the setting of distillation onto a larger student network, showing the episodic returns collected during distillation. Distillation was performed over five random seeds and all environments in each of three suites ‘atari’, ‘mujoco’, and ‘procgen’. The final training performance of the teacher models is shown as a dashed line. Results for procgen are calculated over training levels. We observe that PPD students achieve higher rewards than student-distill ones, and are generally more sample efficient. The teacher-distill trace is omitted here because it reflects performance on teacher-generated trajectories, leading to misleadingly high training scores; the complete curve is shown and discussed as Supplementary Figure 8. Each curve is the interquartile mean over all environments in the suite, with 95% confidence intervals.
@@
We further show that PPD substantially outperforms a baseline with only PPO (without distillation) in Appendix B.4.
Table 1: Performance of student models of three sizes (smaller, same, larger), trained using the three distillation methods (PPD, student-distill, and teacher-distill). Evaluation is performed in a test setting (which in the case of ‘procgen’ corresponds to using a different set of levels). Results are calculated as fraction of teacher score for each environment and random seed, and then aggregated by geometric interquartile mean (95% confidence intervals are included in brackets).
env smaller
td sd ppd
atari 0.86x [0.81, 0.9] 0.94x [0.89, 1.0] 0.97x [0.94, 1.01]
mujoco 0.99x [0.98, 1.0] 0.99x [0.98, 1.0] 0.99x [0.97, 1.0]
procgen   0.72x [0.69, 0.77]    0.78x [0.74, 0.83] 0.93x [0.87, 1.01]
same-size
td sd ppd
atari 1.0x [0.97, 1.04]   1.07x [1.03, 1.12]     1.08x [1.04, 1.13]
mujoco 0.99x [0.98, 1.0] 1.0x [0.99, 1.0] 1.0x [1.0, 1.02]
procgen   0.84x [0.78, 0.89]    0.88x [0.84, 0.94]     1.03x [0.97, 1.08]
larger
td sd ppd
atari   1.07x [1.03, 1.13]    1.09x [1.05, 1.15] 1.1x [1.05, 1.16]
mujoco 0.99x [0.97, 1.0] 1.0x [1.0, 1.02] 1.0x [0.99, 1.01]
procgen   0.91x [0.86, 0.98]   0.96x [0.93, 1.0] 1.04x [0.98, 1.08]
@@
Table 2: Performance of student models, trained using the three distillation methods (PPD, student-distill, and teacher-distill) using ‘imperfect teachers’ that are artificially corrupted to decrease in performance. Results are calculated as a fraction of the original teacher score for each environment and random seed, and then aggregated by geometric interquartile mean (95% confidence intervals are included in brackets). Distillation is performed on four Atari and four Procgen environments, and onto larger student networks.
@@
Figure 2: Performance of student models trained using PPD with different values of $\lambda\in\left\{0.5,1,2,5\right\}$, over 5 random seeds. Distillation is performed on four Atari environments and onto larger student networks. Dashed lines indicate $100\%$ (black) and $90\%$ (gray) of the performance of the teacher models. All curves denote interquartile means with 95% confidence intervals. A circle on each curve marks the last training step where the student’s performance is below 90% of the teacher’s. All curves were collected during training, as in Figure 1.
@@ 4 Related Work
Many current policy distillation methods overlook the rewards obtained by the student during the distillation process. Instead, they frame policy distillation solely as a supervised learning task. In this approach, a dataset of state observations and teacher outputs is collected using either the student or teacher policies. The data is then stored in a distillation dataset from which training mini-batches are sampled.
A significant body of recent work focuses on distillation from teachers with privileged information or full observability to partially-observable students (Nguyen et al., 2022; Weihs et al., 2021), where pure supervised learning cannot fully capture the teacher’s behavior. In these scenarios, the combination of RL with distillation becomes crucial for effective learning. While PPD was primarily designed for fully-observable settings, its integration of environment rewards with distillation makes it potentially suitable for such scenarios, as suggested by its robustness to imperfect teachers (albeit in a different context).
@@
Meanwhile, PD as a form of supervised learning can be easily adapted to transfer functions between different types of models, such as value-based teachers (e.g., trained with DQN) and policy-based students. This proves beneficial in scenarios like reincarnating reinforcement learning (Agarwal et al., 2022), allowing continued training with alternate RL methods after distillation. For example, Green et al. (2019) showed that teachers trained using DQN can be distilled into “PPO students,”, allowing for subsequent fine-tuning with Proximal Policy Optimization (PPO) (Green et al., 2019). However, unlike in our work, PPO was not used during distillation, and the trajectories for the distillation loss were collected in a teacher-driven way, with the teacher interacting with the environment instead of the student. This method closely resembles the ‘teacher_distill’ strategy in our baseline, which we demonstrated to be less effective than PPD.
@@
An alternative way to perform policy distillation is through imitation learning (IL), where demonstrations from teacher models provide specific actions as hard targets, rather than using the typical soft targets (e.g., output probabilities). For example, Zhuang et al. (2023) use DAgger (Dataset Aggregation) (Ross et al., 2011) to distill five behaviors for a robot quadruped (climb, leap, crawl, tilt, run) into a single control policy, by selecting the appropriate teacher for each part of a training obstacle course. The advantage of DAgger over naive Behavior Cloning stems from its ability to query teacher actions for new states encountered by the student, which is always possible in policy distillation, rather than relying solely on teacher-collected demonstrations. The combination of IL with RL has also been explored to improve the sample efficiency of RL methods and to bias behaviors learned with RL towards natural movements (Rajeswaran et al., 2017).
Most relevant to our approach are works that combine policy distillation with reinforcement learning in different ways. COSIL (Nguyen et al., 2022) and TGRL (Shenfeld et al., 2023) incorporate the distillation loss in the reward function, while ADVISOR (Weihs et al., 2021) adds it to a policy gradient objective like in PPD, although without clipping. All three methods, however, focus exclusively on settings with privileged teachers and partially observable students, using different methods to dynamically trade off between teacher guidance and environment rewards. Similarly, (Schmitt et al., 2018) propose using Population Based Training (Jaderberg et al., 2017) to dynamically update this relative weighting. The main difference with Proximal Policy Distillation is that PPD extends PPO by applying proximal constraints to both objectives and enabling effective student-driven exploration, while these approaches rely on balancing schemes and, in the case of COSIL and TGRL, are based on different RL algorithms (SAC and IMPALA respectively). ADVISOR, like PPD, builds on PPO but differs in how it combines the objectives. Additionally, PPD is explicitly evaluated in settings with architectural differences between teacher and student, and shows robustness to imperfect teachers even without dynamic loss balancing. While these prior works focus on the specific challenge of partial observability, we evaluate PPD across a wide range of standard environments including discrete and continuous control tasks, demonstrating its effectiveness as a general policy distillation framework.
@@ 5 Discussion and Conclusions
Future work should focus on extending PPD to use teachers as skills priors, especially in the case of multitask policy distillation. To make that practical, and to further help PPD students to achieve higher performance than their teacher, it will be beneficial to implement a dynamic balancing in the trade-off between the PPO and individual teacher distillation losses, as for example done by Schmitt et al. (2018); Weihs et al. (2021); Nguyen et al. (2022); Shenfeld et al. (2023).